Hi..I suck ASS when it comes to math! So Please help me with the following
-2(-x+8)+3(2x+1)=131
3(x-2)=6
5(x-2)=115
-2x+5-3x=-90
2x-24=34
-3x-18=-x-44
46-3x=-23
3x-16.8=34.2
2x-3.5=16.5
Whoever answer this you the real MVP and need to hmu so you can help me one on one!!



Answer by Alan3354(69443)

-2(-x+8)+3(2x+1)=131
===================
Do the easier ones first.
3(x-2)=6
Divide by 3.
Add 2.
===========
5(x-2)=115
Similar to above.
===============
-2x+5-3x=-90
2x-24=34
Add 24
Divide by 2.
==============
-3x-18=-x-44
Similar
=============
46-3x=-23
Similar
============
3x-16.8 = 34.2
Add, then divide.
==========
2x-3.5=16.5
More of the same.
=======================
email via the TY note to check your work, or for assistance.
---
idk what hmu means.